OpenStudy (anonymous):

Find the value of constants c and d that make the function below continuous at x = 2. x^2-3x x<2 F(x) c x=2

OpenStudy (anonymous):

whoops...and d+x x>2 is the last part

OpenStudy (anonymous):

For it to be continuous, the values of each function would have to be equal at 2. In mathematical terms: \[x^2-3x = c = d+x\]

OpenStudy (anonymous):

so like x^2-3x=d+x=2?

OpenStudy (anonymous):

They don't equal 2 I'm afraid. Take the first expression for example: \(f(x) = x^2-3x\) What would f(x) equal at x=2?

OpenStudy (anonymous):

no it would at 3 though?

OpenStudy (anonymous):

I mean what is f(2)?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

Correct! Now going back to the equation I wrote earlier: \[x^2-3x = c = d+x \text{ ; } x=2\]You should be able to work out the values of c and d.